OpenStudy (anonymous):

What is the value of the rational expression below when x is equal to 3?

OpenStudy (anonymous):

17 +x /x+2

OpenStudy (anonymous):

The answer is 4.

OpenStudy (whpalmer4):

Assuming you mean \[\frac{17+x}{x+2}=\frac{17+3}{3+2}=\frac{20}{5}=4\]

OpenStudy (anonymous):

so how does that work?

OpenStudy (anonymous):

Huh? you simply substitute for x. That's all how it works..

OpenStudy (whpalmer4):

How does what work? I just replaced \(x\) with 3 wherever I encountered it, then did the resulting arithmetic.